OpenStudy (anonymous):

Let z = 8e^(5iπ/6) Write z^6 in the form of a+bi this is what I got, 8^6(cos(20π/3)+isin(20π/3)) is it correct?

OpenStudy (anonymous):

no i don't think so

OpenStudy (anonymous):

the \(8^6\) part is right, your last job is to multiply the angles by 6

OpenStudy (anonymous):

\[z = 8e^{\frac{5iπ}{6}\] \[z^6=( 8e^{\frac{5iπ}{6})^6\] \[=8^6e^{5p\pi}\]

OpenStudy (valpey):

\[z = 8e^{\frac{5i\pi}{6}}\] \[z^{6} = 8e^{5i\pi} = 8(e^{i\pi})^{5}\] \[ = 8(-1)^{5} = -8\]

OpenStudy (anonymous):

\[z = 8e^{\frac{5iπ}{6}}\]

OpenStudy (anonymous):

oh oops i found my mistake, I was multiplying the angle by 8..

OpenStudy (anonymous):

\[z^6 = (8e^{\frac{5iπ}{6}}^6=8^6e^{5\pi i}\]

OpenStudy (valpey):

WE MUSN'T FORGET \[e^{i\pi} = -1 \] (therefore god exists)

OpenStudy (anonymous):

in anycase even though i cannot see what i am writing, you should get \[8^6(\cos(5\pi)+i\sin(5\pi))\]

OpenStudy (valpey):

Oh, wait, I probably mean \[-8^{6} = -2^{18}\]

OpenStudy (anonymous):

man is this slow \(-8^6\) is your answer

OpenStudy (anonymous):

yep that's what I got thanks :)